Show that the period of oscillation on a main helical spring is given by T=2π√{k/m}

1
Expert's answer
2022-03-10T10:17:59-0500

Answer:

According to Newton's second law we have: "a=\dfrac{F}{m}=-\dfrac{k}{m}x;"

"a=-w^2x;"

Equating the two formulas for acceleration we obtain:

"-w^2x=-\dfrac{k}{m}x; w^2=\dfrac{k}{m}; w=\sqrt{\dfrac{k}{m}};"

"T=\dfrac{2\pi}{\omega}; T=2\pi\sqrt{\dfrac{m}{k}}."